A 1000-g ball is tied to a string 60cm long and is whirled in a vertical circle at 1 revolution per second. if the tension at the highest point is zero what is the speed of the body at the highest point?

1
Expert's answer
2022-01-12T08:28:21-0500

The second Newton's law for the highest point:


"ma = mg+T"

where "a = \dfrac{v^2}{r}" is the centripetal acceleration ("v" is the speed of the ball, "r = 60cm = 0.6m" is the radius of the circle), and "T = 0" is the tension of the string. "m = 1000g = 1kg" and "g = 9.81m\/s^2" is the gravitational acceleration. Thus, expressing "v" obtain:


"m\dfrac{v^2}{r} = mg\\
v = \sqrt{rg}\\
v = \sqrt{0.6\cdot 9.81} \approx 2.4m/s"

Answer. 2.4m/s.